OpenStudy (anonymous):

Quick question what is (-|x|)^2 equal out to?

OpenStudy (blockcolder):

\[ (-|x|)^2=|x|^2=|x^2|=x^2\] since x^2>=0 for all real x.

OpenStudy (anonymous):

cool thanks